OpenStudy (anonymous):

Lecture 5 - Parametric equations Hi, When the cycloid is first introduced, the position of a point, P, on the circumference is described as (all the following are vectors) OP = OA + BA + BP BP = < -asinTheta, -acosTheta > I don't understand this as I thought x = rcosTheta y = rsinTheta Therefore should the components of BP be, BP = < -acosTheta, -asinTheta >? Thanks

OpenStudy (phi):

yes, for a circle centered at the origin, and measuring theta counter-clockwise from the x-axis, then the points on it circumference have coords \[ x= r \cos \theta \ , y= r \sin \theta \] but in this problem, the angle is measured in a "non-standard" way. specificially